A proton (mass m = 1.67 × 10-27 kg) is being accelerated along a straight line at 8.70 × 1011 m/s2 in a machine. If the proton has an initial speed of 3.20 × 106 m/s and travels 1.30 cm, what then is (a) its speed and (b) the increase in its kinetic energy?
Answer #1

a) using 22-u? = 2 as 202 = u2t2a5 202 = (3.2 * 105m/s)2 + 2 (8.7 *10^-/s2X0.013m) Wat e = 3.2035 * 10% m/s c DKE = 4 m/122-4